OpenStudy (dimples97):

What is the simplified form of this expression? Could someone teach me how to answer it if my answer is wrong? This is the problem

OpenStudy (dimples97):



OpenStudy (dimples97):

this is my answer x9

zepdrix (zepdrix):

Use the caret symbol when you want to show an exponent in text. So \(\rm x^9\) you would write this way x^9.

zepdrix (zepdrix):

So you're saying that \(\large\rm x^{-9}\) turned into \(\rm x^9\) ? Hmmm

zepdrix (zepdrix):

When dealing with negative exponents, it can be helpful to rewrite your value as a fraction. Remember that you can write any number over 1 to turn it into a fraction, \(\large\rm \dfrac{x^{-9}}{1}\) We can change our exponent to positive by throwing the whole thing into the denominator,\[\large\rm \frac{}{1x^9}\]Hmm but what's left upstairs? :O Nothing? A 1 is still there actually!\[\large\rm \frac{1}{1x^9}\]Which we can simply write as,\[\large\rm \frac{1}{x^9}\]I hope the extra 1's and fraction business didn't make things too confusing. Exponents can be a little tricky.

zepdrix (zepdrix):

What do you think Dimps? :d

OpenStudy (dimples97):

Ohhh okay i see i see

OpenStudy (dimples97):

Do you mind helping me with one more zep ? D: just so I'm sure ya know?

zepdrix (zepdrix):

sure

OpenStudy (dimples97):